OpenStudy (anonymous):

a college bookstore ordered six boxes of red pens. the store sold 38 red pens last week and 49 red pens this week. 9 pens were left on the shelf. how many pens in each box? it this how I solve this...? x-38 * 49=9

OpenStudy (callisto):

the number of pens in each box = (38+49+9)/6 = 16

OpenStudy (callisto):

or let x be the no. of pens in each box 6x = 38+49+9 you can get the same answer

OpenStudy (anonymous):

oh ok that makes sense cuz i kept doin the problem & kept gettin it wrong thank youfor ur help

OpenStudy (anonymous):

Neenee I'm working on the same kind of problems